# MS-Mix: Unveiling the Power of Mixup for Multimodal Sentiment Analysis

### v1

@@ 3. Methodology
In this section, we first introduce the task definition and the overview of our MS-Mix framework. We then describe the details of each module in MS-Mix.
3.1. Task Definition and Framework Overview
@@
Given a MSA dataset with $n$ samples, $\textbf{X}=(\textbf{x}_{1},\textbf{x}_{2},...,\textbf{x}_{n})$} and $\textbf{Y}=(y_{1},y_{2},...,y_{n})$} are the data and labels. For the $i_{th}$ sample $\textbf{x}_{i}\in\textbf{X}$, where $\textbf{x}_{i}$ comprises data from the text (t), video (v), and audio (a) modalities, denoted as $\{\textbf{x}^{t}_{i},\textbf{x}^{v}_{i},\textbf{x}^{a}_{i}\}$. The goal of the MSA task is to learn a mapping function $\mathcal{F}:\textbf{X}\mapsto\textbf{Y}$ to predict the occurrence of each emotion category.
@@
MS-Mix is a novel data augmentation method in the latent space, acting on the multimodal features output by the encoder, and is applied before feature fusion. As illustrated in Fig.2, the input to MS-Mix consists of multimodal features derived from the model’s encoder outputs. The SASS strategy first identifies feature pairs suitable for mixing, which are then adaptively blended through the SIG module to generate augmented features. Both the original and synthesized features are subsequently fed into the fusion module. Additionally, the SAL serves as an auxiliary regularization term to jointly optimize the entire network and the SIG module.
@@
As a common starting point for many mixup-based augmentation methods, vanilla mixup (Zhang et al., 2018) employs a ratio $\lambda$ to construct a mixed sample ($\hat{\textbf{x}},\hat{y}$) by performing global linear interpolation directly on the sample pair $\{(\textbf{x}_{i},y_{i}),(\textbf{x}_{j},y_{j})\}$:
(1) $\hat{\textbf{x}}=\lambda\cdot\textbf{x}_{i}+(1-\lambda)\cdot\textbf{x}_{j},$
@@
Mixing of original data can easily lead to sentiment semantic confusion in generated data (Verma et al., 2019). Therefore, for each batch of data, we choose to learn different mixing ratios $\lambda^{m}$ to mix the features $\textbf{Z}^{m}=f_{\theta}(\textbf{X})\in\mathbb{R}^{B\times d^{m}}$ of each modal in the latent space separately (Eq. 2), and calculate the label ratio $\lambda^{L}$ to mix the labels as follows:
(2) $\hat{\textbf{z}}^{m}=\mathrm{H}(f_{\theta}(\textbf{x}_{i}^{m}),f_{\theta}(\textbf{x}_{j}^{m}),\lambda^{m}),\textbf{x}^{m}\in\textbf{X},$
(3) $\hat{y}=\mathrm{G}((y_{i},y_{j}),\bm{\lambda}^{L}),y\in\textbf{Y}.$
where $\hat{\textbf{z}}^{m}$ represents the the mixed features, $m\in\{t,v,a\}$ represents a modality belonging to the set of the three modalities in the dataset, $B$ is the mini-batch size, $d^{m}$ is the hidden space dimension of modality $m$, and $f_{\theta}$ is the backbone.
Finally, the mixed features and the original features will be fused into a whole set for the subsequent feature fusion process.
3.2. Sentiment-Aware Sample Selection
Emotional expressions exhibit significant variation across samples. While random in-batch mixing enhances the feature smoothness, it often ignores semantic consistency in the emotional space. This may lead to the mixing of two semantically opposite samples with strong emotion into a mixed sample with a neutral label and semantic confusion features. Therefore, we proposed the SASS strategy based on the emotional semantic distance, and screened the feature samples within a batch before mixing. Specifically, we use the feature cosine distance between each sample to represent the semantic information similarity (Pennington et al., 2014), and mix the sample pairs with a similarity greater than the threshold $\delta$ within each batch. This ensures that only samples with analogous emotions are combined, a critical factor overlooked by previous methods (Georgiou et al., 2024).
@@ 3.2. Sentiment-Aware Sample Selection
(4) $\mathbf{Z}^{m}_{\text{norm}}=\frac{\mathbf{Z}^{m}}{\sqrt{\sum_{i=1}^{d^{m}}(Z_{i}^{m})^{2}}}.$
@@
(5) $\textbf{S}=\sum_{m\in t,v,a}{\mathbf{Z}^{m}_{\text{norm}}}\cdot(\mathbf{Z}^{m}_{\text{norm}})^{\top}/3.$
@@
Since the matrix S is symmetric ($s_{ij}=s_{ji},s\in\textbf{S}$) and the diagonal elements represent the self-similarity, we define the normalized upper triangular matrix $\textbf{S}^{\prime}$ as:
(6) $\textbf{S}^{\prime}=\frac{1}{n(n-1)}\sum_{i\neq j}s_{ij}.$
@@
Then we randomly selected $B$ pairs to mix from the feature pairs $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\},\textbf{z}\in\textbf{Z}$, with a similarity greater than $\delta$ = 0.2. Our SASS strategy employs the similarity threshold to exclude sample pairs with opposite emotions before mixing. A lower threshold effectively prevents semantic confusion in the mixed samples while maximizing data diversity.
3.3. Sentiment Intensity Guided Mixing Module
@@
The mix ratio $\lambda$ is one of the most important hyperparameters in mixup-based methods, used to balance the degree of mixing between two or more samples. $\lambda$ is usually sampled from the $Beta(\alpha,\alpha)$ distribution. However, in recent years, some methods (Kim et al., 2020b; Liu et al., 2022; Qin et al., 2024a) have been able to automatically optimize the mixture proportions based on the sample states, thereby achieving better matching between the data and the label.
@@
To ensure that samples with richer emotional semantics contribute more substantially to the mixing process, we propose the SIG mixing module. Specifically, for each modality, a Multi-Head self-Attention (MHA) (Vaswani et al., 2017) encoder is trained to predict emotional intensity values $\textbf{I}^{m}$. This encoder comprises an MHA layer, a residual connection, layer normalization, and a tanh activation function. The modality-specific emotional intensity predictor is denoted as $f_{\phi}^{m}$ (Eq. 7-9). These predictions are then utilized as weighting factors to adjust the mixing ratios $\lambda^{m}$ during augmentation dynamically.
@@
(7) $head_{i}=\mathrm{Softmax}\left(\frac{\textbf{Z}^{m}\textbf{W}_{i}^{Q}\cdot\textbf{Z}^{m}({\textbf{W}_{i}^{K}})^{\top}}{\sqrt{d_{k}}}\right)\textbf{Z}^{m}\textbf{W}_{i}^{V},$
(8) $\mathrm{MHA}(\textbf{Z}^{m})=\mathrm{LN}(\mathrm{Concat}(head_{1},head_{2},.,head_{h})\textbf{W}^{O}+\textbf{Z}^{m}),$
(9) $\textbf{I}^{m}=f_{\phi}^{m}(\textbf{Z}^{m})=\mathrm{tanh}(\mathrm{GobalPool}(\mathrm{MHA}(\textbf{Z}^{m}))).$
@@
where $\textbf{W}^{\{Q,K,V\}}={\{\textbf{W}^{Q},\textbf{W}^{K},\textbf{W}^{V}\}}$ are learnable weight matrices that map the input features to queries $Q$, keys $K$, and values $V$, $d_{k}$ is the dimension of $K$, LNLN denotes Layer Normalization (Ba et al., 2016), the number of attention heads $h$ is set to 4 or 6, and $\textbf{W}^{O}$ is the output linear transformation matrix.
@@
Based on the intensity of emotions $\textbf{I}^{m}$ in Eq. 9, we perform Min-Max normalization on it, which can further calculate the intermediate mixing weights ${\bm{\omega}}^{m}_{i}$ for each feature pair $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\}$:
(10) $\bm{\omega}_{i}^{m}=\frac{|\textbf{I}_{i}^{m}|-min(\textbf{I}^{m})}{max(\textbf{I}^{m})-min(\textbf{I}^{m})+\epsilon}.$
@@
(11) $\lambda^{m}_{i,j}=(\frac{\bm{\omega}_{i}^{m}}{\bm{\omega}_{i}^{m}+\bm{\omega}_{j}^{m}}+\lambda_{base})/2,$
@@
(12) $\lambda^{L}_{i,j}=\sum_{m\in t,v,a}\lambda_{i,j}^{m}/3.$
Finally, we can get the mixed feature $\hat{\textbf{z}}^{m}$ and mixed label $\hat{y}$ using Eq. 2-3. Unlike the previous Mixup-based studies (Zhang et al., 2018; Verma et al., 2019; Yun et al., 2019; Uddin et al., 2021), our proposed SIG module adaptively determines the mixing ratios between samples based on sentiment intensity, maintaining end-to-end trainability and allowing for continuous optimization throughout the training process.
3.4. Sentiment Alignment Loss Function
To enhance the accuracy of the emotion intensity predictor, we introduce the SAL as an additional regularization term to align the predicted sentiment distribution with the ground-truth labels. In our task, we seek the mapping from the data $x$ to its label $y$ modeled by network $f_{\theta}$, where the network parameters $\theta$ are optimized by minimizing a loss function $\mathcal{L}_{task}$ using Mean Squared Error (MSE):
(13) $\min_{\theta}\mathcal{L}_{task}(f_{\theta}(\textbf{x}),y).$
Since the emotional labels are continuous values (Zadeh et al., 2016, 2018b; Yu et al., 2020), we consider the mixup regression task. Therefore, we get the mapping between the mixed $\hat{\textbf{x}}$ and $\hat{y}$ by optimizing mixed MSE loss $\mathcal{L}_{mixMSE}$ (Eq.14). By minimizing $\mathcal{L}_{mixMSE}$ (Eq.15), the consistency between the mixed samples and labels is constrained.
(14) $\mathcal{L}_{mixMSE}=\lambda^{L}_{i,j}\cdot MSE(f_{\theta}(\hat{\textbf{x}_{i}}),y_{i})+(1-\lambda^{L}_{i,j})\cdot MSE(f_{\theta}(\hat{\textbf{x}_{j}}),y_{j}).$
@@ 3.4. Sentiment Alignment Loss Function
The KL divergence (Van Erven and Harremos, 2014) is commonly used to measure the amount of information lost when an approximate distribution is used to represent a true distribution. A smaller KL divergence indicates that the two distributions are more similar. Thus, in this study, we introduce the SAL based on the KL divergence to align the predicted sentiment distribution with the ground-truth labels, thereby enhancing the accuracy of the emotion intensity predictor. Specifically, we convert $\textbf{I}^{m}$ and ground truth Y into probability distributions $\textbf{P}^{m}$ and $\textbf{P}^{L}$ (Eq.16), then calculate the KL divergence between them (Eq.17). Finally, we scale the KL divergence and incorporate it as an additional regularization term (Eq.18). The scale factor $\beta$ is $10^{3}$. This approach further promotes alignment across different modalities in emotion label prediction.
(16) $\textbf{P}^{m}=\frac{exp(\textbf{I}^{m})}{\sum^{B}_{i=1}exp(\textbf{I}^{m}_{i})},\textbf{P}^{L}=\frac{exp(\textbf{Y})}{\sum^{B}_{i=1}exp(y_{i})},$
(17) $\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})=\frac{1}{B}\sum^{B}_{i=1}\textbf{P}^{L}_{i}(log\textbf{P}^{L}_{i}-log\textbf{P}^{m}_{i}),$
(18) $\mathcal{L}_{SAL}=\sum_{m\in{t,v,a}}\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})\cdot\beta.$
@@
(19) $\mathcal{L}_{total}=\mathcal{L}_{task}+\xi_{1}\cdot\mathcal{L}_{mixMSE}+\xi_{2}\cdot\mathcal{L}_{SAL}.$
where $\xi_{1}$ and $\xi_{2}$ are the weight hyperparameters used to regulate mixed MSE loss and SAL loss, respectively, with the effects of $\xi_{1}$ and $\xi_{2}$ to be evaluated in our ablation study. The SAL can serve as an auxiliary regularization term that not only facilitates the training of the modality-specific predictor $f_{\phi}^{m}$, but also promotes consistency across different modalities.
4. Experiments
To evaluate the proposed approach, we conducted comprehensive experiments on three public benchmark datasets for MSA: CMU-MOSI (MOSI) (Zadeh et al., 2016), CMU-MOSEI (MOSEI) (Zadeh et al., 2018b), and CH-SIMS (SIMS) (Yu et al., 2020). We compare our method against three representative Mixup-based methods (Verma et al., 2019; Venkataramanan et al., 2023; Georgiou et al., 2024) on six state-of-the-art backbones (Zadeh et al., 2017; Liu and Shen, 2018; Tsai et al., 2019; Hazarika et al., 2020; Zhang et al., 2023; Zhuang et al., 2024).
4.1. Datasets

### v2

@@ 3 Methodology
In this section, we first introduce the task definition and the overview of our MS-Mix framework. We then describe the details of each module in MS-Mix. The complete MS-Mix process is presented in Algorithm 1.
Input: Data X and label Y, backbone $f_{\theta}$, prediction head $f_{\psi}$,
hyperparameters $\alpha$, $\delta$, $\beta$
1
2Initialize parameters;
3
4for each epoch do
5   Sample mini-batch $B$;
6
7 // Step 1: Sentiment-Aware Sample Selection Extract features $\textbf{Z}^{m}=f_{\theta}(\textbf{X})$ for each modality;
8   Compute averaged similarity across modalities;
9   Select consistent pairs where $s_{ij}>\delta$;
10
11 for each consistent pair $((x_{i},y_{i}),(x_{j},y_{j}))$ do
12    // Step 2: Intensity-Guided Mixing
13      Predict intensities $I^{m}$ via $\textbf{Z}^{m}$;
14      Obtain modality specific mixing ratios $\lambda^{m}_{i,j}$, and mixing ratios $\lambda^{L}_{i,j}$ of labels;
15
16     Generate mixed features $\hat{\textbf{z}}^{m}=\lambda^{m}_{i,j}\textbf{z}_{i}^{m}+(1-\lambda^{m}_{i,j})\textbf{z}_{j}^{m}$, and mixed label $\hat{y}=\lambda^{L}_{i,j}y_{i}+(1-\lambda^{L}_{i,j})y_{j}$ ;
17
18    // Step 3. Losses & Update
[…]
21      Alignment loss: $\mathcal{L}_{\text{SAL}}\leftarrow\text{KL}(P^{L}||P^{m})$;
22      Total loss: $\mathcal{L}_{total}=\mathcal{L}_{task}+\xi_{1}\cdot\mathcal{L}_{mixMSE}+\xi_{2}\cdot\mathcal{L}_{SAL}$;
23
24     Update $\theta$ by minimizing $\mathcal{L}_{total}$;
25
26   end for
27
28 end for
29return Trained $f_{\psi}\circ f_{\theta}$;
Algorithm 1 Training Procedure of MS-Mix
3.1 Task Definition and Framework Overview
@@
Given a MSA dataset with $n$ samples, $\textbf{X}=(\textbf{x}_{1},\textbf{x}_{2},...,\textbf{x}_{n})$} and $\textbf{Y}=(y_{1},y_{2},...,y_{n})$} where $y_{i}$ are continuous sentiment intensity scores, each $\mathbf{x}_{i}\in\mathbf{X}$ comprises data from the text ($t$), video ($v$), and audio ($a$) modalities, denoted as $\{\mathbf{x}_{i}^{t},\mathbf{x}_{i}^{v},\mathbf{x}_{i}^{a}\}$. The goal of the MSA task is to learn a mapping function $\mathcal{F}:\mathbf{X}\mapsto\mathbf{Y}$ to predict the continuous intensity of each emotion category. This mapping is realized by a backbone $f_{\theta}$ followed by a prediction head $f_{\psi}$.
@@
To address data scarcity and enhance generalization in this regression-based MSA task, we propose MS-Mix. This novel latent-space data augmentation method operates on the multimodal features output by the encoder and is applied before feature fusion. As illustrated in Fig.2, the input to MS-Mix consists of multimodal features derived from the model’s encoder outputs. The SASS strategy first identifies feature pairs suitable for mixing, which are then adaptively blended through the SIG module to generate augmented features. Both the original and synthesized features are subsequently fed into the fusion module. Additionally, the SAL serves as an auxiliary regularization term to jointly optimize the entire network and the SIG module.
@@
As a common starting point for many mixup-based augmentation methods, vanilla mixup [13] employs a ratio $\lambda$ to construct a mixed sample ($\hat{\textbf{x}},\hat{y}$) by performing global linear interpolation directly on the sample pair $\{(\textbf{x}_{i},y_{i}),(\textbf{x}_{j},y_{j})\}$:
$\hat{\textbf{x}}=\lambda\cdot\textbf{x}_{i}+(1-\lambda)\cdot\textbf{x}_{j},$ (1)
@@
Mixing of original data can easily lead to sentiment semantic confusion in generated data [34]. Therefore, for each batch of data, we choose to learn different mixing ratios $\lambda^{m}$ to mix the features $\textbf{Z}^{m}=f_{\theta}(\textbf{X})\in\mathbb{R}^{B\times d^{m}}$ of each modal in the latent space separately (Eq. 2), and calculate the label ratio $\lambda^{L}$ to mix the labels as follows:
$\hat{\textbf{z}}^{m}=\mathrm{H}(f_{\theta}(\textbf{x}_{i}^{m}),f_{\theta}(\textbf{x}_{j}^{m}),\lambda^{m}),\textbf{x}^{m}\in\textbf{X},$ (2)
$\hat{y}=\mathrm{G}((y_{i},y_{j}),\bm{\lambda}^{L}),y\in\textbf{Y},$
where $\hat{\textbf{z}}^{m}$ represents the mixed features, $m\in\{t,v,a\}$ represents the set of the three modalities in the dataset, $B$ is the mini-batch size, $d^{m}$ is the hidden space dimension of modality $m$, and $f_{\theta}$ is the backbone.
Finally, the mixed and original features will be fused into a single set for subsequent feature fusion.
3.2 Sentiment-Aware Sample Selection
Emotional expressions exhibit significant variation across samples. While random in-batch mixing enhances feature smoothness, it often fails to preserve semantic consistency in the emotional space. This may lead to the mixing of two semantically opposite samples with strong emotion into a mixed sample with a neutral label and semantic confusion features. Therefore, we proposed the SASS strategy based on emotional semantic distance and screened the feature samples within each batch before mixing. Specifically, we use the cosine distance between each sample to represent semantic information similarity [38], and mix sample pairs with similarity greater than the threshold $\delta$ within each batch. This ensures that only samples with analogous emotions are combined, a critical factor that previous methods overlooked [17].
@@ 3.2 Sentiment-Aware Sample Selection
$\mathbf{Z}^{m}_{\text{norm}}=\frac{\mathbf{Z}^{m}}{\sqrt{\sum_{i=1}^{d^{m}}(Z_{i}^{m})^{2}}}.$ (3)
@@
$\textbf{S}=\sum_{m\in t,v,a}{\mathbf{Z}^{m}_{\text{norm}}}\cdot(\mathbf{Z}^{m}_{\text{norm}})^{\top}/3.$ (4)
@@
Since the matrix S is symmetric ($s\in\textbf{S},s_{ij}=s_{ji}$) and the diagonal elements represent the self-similarity, we define the normalized upper triangular matrix $\textbf{S}^{\prime}$ as:
$\textbf{S}^{\prime}=\frac{1}{n(n-1)}\sum_{i\neq j}s_{ij}.$ (5)
@@
We randomly selected $B$ pairs to mix from the feature pairs $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\},\textbf{z}\in\textbf{Z}$, with a similarity greater than $\delta$ = 0.2. Our SASS strategy employs the similarity threshold to exclude sample pairs with opposite emotions before mixing. A lower threshold effectively prevents semantic confusion in the mixed samples while maximizing data diversity.
3.3 Sentiment Intensity Guided Mixing Module
@@
The mixing ratio $\lambda$ is one of the most important hyperparameters in mixup-based methods, controlling the degree of mixing between two or more samples. $\lambda$ is usually sampled from the $Beta(\alpha,\alpha)$ distribution. However, in recent years, some methods [39, 30, 33] have automatically optimized mix ratios based on sample states, thereby achieving better alignment between the data and the labels.
@@
To ensure that samples with richer emotional semantics contribute more substantially to the mixing process, we propose the SIG mixing module. Specifically, for each modality, a Multi-Head self-Attention (MHA) [40] encoder is trained to predict emotional intensity values $\textbf{I}^{m}$. This encoder comprises an MHA layer, a residual connection, layer normalization, and a tanh activation function. The modality-specific emotional intensity predictor is denoted as $f_{\phi}^{m}$ (Eq. 6-8). These predictions are then used as weighting factors to adjust the mixing ratios $\lambda^{m}$ during augmentation dynamically.
@@
$head_{i}=\mathrm{Softmax}\left(\frac{\textbf{Z}^{m}\textbf{W}_{i}^{Q}\cdot\textbf{Z}^{m}({\textbf{W}_{i}^{K}})^{\top}}{\sqrt{d_{k}}}\right)\textbf{Z}^{m}\textbf{W}_{i}^{V},$ (6)
$\mathrm{MHA}(\textbf{Z}^{m})$ $=\mathrm{LN}(\mathrm{Concat}(head_{1},head_{2},.,head_{h})\textbf{W}^{O}+\textbf{Z}^{m}),$ (7)
$\textbf{I}^{m}=f_{\phi}^{m}(\textbf{Z}^{m})=\mathrm{tanh}(\mathrm{GobalPool}(\mathrm{MHA}(\textbf{Z}^{m}))).$ (8)
@@
where $\textbf{W}^{\{Q,K,V\}}={\{\textbf{W}^{Q},\textbf{W}^{K},\textbf{W}^{V}\}}$ are learnable weight matrices that map the input features to queries $Q$, keys $K$, and values $V$, $d_{k}$ is the dimension of $K$, LNLN denotes Layer Normalization [41], the number of attention heads $h$ is set to 4 or 6, and $\textbf{W}^{O}$ is the output linear transformation matrix.
@@
Based on the intensity of emotions $\textbf{I}^{m}$ in Eq. 8, we perform Min-Max normalization on it in modality, which can further calculate the intermediate mixing weights ${\bm{\omega}}^{m}_{i}$ for each feature pair $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\}$:
$\bm{\omega}_{i}^{m}=\frac{|\textbf{I}_{i}^{m}|-min(\textbf{I}^{m})}{max(\textbf{I}^{m})-min(\textbf{I}^{m})+\epsilon},$ (9)
@@
$\lambda^{m}_{i,j}=(\frac{\bm{\omega}_{i}^{m}}{\bm{\omega}_{i}^{m}+\bm{\omega}_{j}^{m}}+\lambda_{base})/2,$ (10)
@@
$\lambda^{L}_{i,j}=\sum_{m\in t,v,a}\lambda_{i,j}^{m}/3.$ (11)
Finally, we can get the mixed feature $\hat{\textbf{z}}^{m}$ and mixed label $\hat{y}$ using Eq. 2. Unlike previous Mixup-based studies [13, 34, 15, 16], our proposed SIG module adaptively determines the mixing ratios between samples based on sentiment intensity, thereby maintaining end-to-end trainability and enabling continuous optimization throughout training.
3.4 Sentiment Alignment Loss Function
To enhance the accuracy of the emotion intensity predictor, we introduce the SAL as an additional regularization term to align the predicted distribution with the ground-truth labels. Formally, our goal is to learn a mapping from an input sample x to its corresponding label $y$ using a network $f_{\psi}\circ f_{\theta}$. The network parameters $\psi$ and $\theta$ are typically optimized by minimizing a task-specific loss $\mathcal{L}_{task}$:
$\min_{\psi,\theta}\mathcal{L}_{task}(f_{\psi}(f_{\theta}(\textbf{x})),y).$ (12)
Since the emotional labels are continuous values [42, 43, 44], we consider the mixup regression task. Therefore, we get the mapping between the mixed $\hat{\textbf{x}}$ and $\hat{y}$ by optimizing mixed Mean Squared Error (MSE) loss $\mathcal{L}_{mixMSE}$ (Eq.13). By minimizing $\mathcal{L}_{mixMSE}$ (Eq.14), the consistency between the mixed samples and labels is constrained.
$\mathcal{L}_{mixMSE}$ $=\lambda^{L}_{i,j}\cdot MSE(f_{\psi}(f_{\theta}(\hat{\textbf{x}_{i}})),y_{i})+(1-\lambda^{L}_{i,j})\cdot MSE(f_{\psi}(f_{\theta}(\hat{\textbf{x}_{j}})),y_{j}).$ (13)
@@ 3.4 Sentiment Alignment Loss Function
The KL divergence [45] is commonly used to measure the amount of information lost when an approximate distribution is used to represent a true distribution. A smaller KL divergence indicates that the two distributions are more similar. Thus, in this study, we introduce the SAL based on the KL divergence to align the predicted sentiment distribution with the ground-truth labels, thereby enhancing the accuracy of the emotion intensity predictor. Specifically, we convert $\textbf{I}^{m}$ and ground truth Y into probability distributions $\textbf{P}^{m}$ and $\textbf{P}^{L}$ (Eq.15), then calculate the KL divergence between them (Eq.16). Finally, we scale the KL divergence and incorporate it as an additional regularization term (Eq. 17). The scale factor $\beta$ is $10^{3}$. This approach further promotes alignment across different modalities in emotion label prediction.
$\textbf{P}^{m}_{i}=\frac{exp(\textbf{I}^{m})}{\sum^{B}_{i=1}exp(\textbf{I}^{m}_{i})},\textbf{P}^{L}_{i}=\frac{exp(\textbf{Y})}{\sum^{B}_{i=1}exp(y_{i})},$ (15)
$\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})=\frac{1}{B}\sum^{B}_{i=1}\textbf{P}^{L}_{i}(log\textbf{P}^{L}_{i}-log\textbf{P}^{m}_{i}),$ (16)
$\mathcal{L}_{SAL}=\sum_{m\in{t,v,a}}\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})\cdot\beta.$ (17)
@@
$\mathcal{L}_{total}=\mathcal{L}_{task}+\xi_{1}\cdot\mathcal{L}_{mixMSE}+\xi_{2}\cdot\mathcal{L}_{SAL}.$ (18)
where $\xi_{1}$ and $\xi_{2}$ are the weight hyperparameters used to regulate mixed MSE loss and SAL loss, respectively, with the effects to be evaluated in Section 4.6. The SAL can serve as an auxiliary regularization term that not only facilitates training of the modality-specific predictor $f_{\phi}^{m}$ but also promotes consistency across modalities.
Figure 3: The t-SNE visualization of the original features and mixed features generated by MS-Mix (a) and $\mathcal{P}$owMix (b) on the MOSEI dataset using the MISA model. We employ a color scheme (blue/red) to differentiate the positive and negative categories, and use transparency to distinguish the original features from the mixed ones.
4 Experiments
To evaluate the proposed approach, we conducted comprehensive experiments on three public MSA datasets: CMU-MOSI (MOSI) [42], CMU-MOSEI (MOSEI) [43], and CH-SIMS (SIMS) [44]. We compare our method against four representative Mixup-based methods [34, 18, 17, 33] on six state-of-the-art backbones [21, 22, 12, 46, 5, 6]. The comparative results are summarized in Tables 1,2,3. In addition, we conducted extensive ablation experiments (Table 4) and parameter sensitivity analysis (Fig. 5), as well as visualizations of the model performance (Fig. 3 and 4).
4.1 Datasets